Ram's age is 4 years less than half of his mother age. Ram's age is 2 years less than thrice the age of her younger brother. If Ram is 13 years old , than find the age of his brother & that of his mother.
Answer by macston(5194)

.
R=Ram's age=13; M=Mother's age; B=Brother's age
.
R=(1/2)M-4
2R=M-4
2R+4=M
2(13)+4=M
26+4=M
30=M
ANSWER 1: Mother is 30 years old.
.
R=3B-2
R+2=3B
(R+2)/3=B
(13+2)/3=B
15/3=B
5=B
ANSWER 2: Brother is 5 years old.